Week of 050808

Open Actions from last week:
  • testing of smoke tests (James - R/GMA - FTS) DONE (R-GMA)
  • new FTS release / deployment plans.
  • Maarten: Look into GRIDVIEW numbers - Phool DONE
  • GRIDVIEW DB to be upgraded (Monday) DONE
  • Check for "lost" files ( Vlado) DONE
  • FTS Sensors - Scripts to write (Gavin) DONE

On shift : James/Patricia

Monday:

Log: Nothing

New Actions:

Discussion:

  • Want to test new FTS1.3 hard against a site DONE
  • gridview update - stop producers for duration DONE
  • Code freeze date for FTS with srmCopy 15th August
  • Got a spreadsheet for resource requirements from CMS for SC3.

  • FIO suggest to not use Phedex pools, but re-use some of the other pools instead. Need to discuss what longer term (service phase) plans for the pools are - 1 or 2 (or 4???) and are they in WAN area or LAN area?

Tuesday:

Log: No

New Actions:

Discussion:

  • R-GMA team will come and give a tutorial when we want one
  • boxes are ready for DPM deployment. VLADO
  • early next week for deployment of 1.3 - working on docs. =

  • Simone: 2.6.0 is out now. Probably no chance of it being installed by Sept 1st. Is this a problem? Probably not...

Wednesday

Log: Nothing to report for service. Tests ongoing with DESY/INFN. INFN seems to be better tuned - 5 streams/10 concurrent.

Actions:

Discussion:

  • How to use the phedex pools for CMS. They're now clean (no replicas).

Thursday

Log: Load noticed on GRID4 - GRIDVIEW from last night. Need to investigate.

Discussion:

  • Found problem with FTS not filling queues -was due to a query to the DB - worked around with a query hint. QF should be released today.
  • Added UNOSAT as another VO for LFC
  • Met FIO/FS yesterday about requirements for Grid machines going to be managed by FIO. Need to state the reliability and storage requirements.

Actions:

Friday

Log: Nothing to report. Rate overnight to INFN/NDGF dropped ~80% with no changes.

Actions:

  • To look at FTS for preasons for dropping rate (Paolo/Gavin)
  • Upgrade UI Node to 2.6

Discussion:

  • Phool found problem with GRIDVIEW DB - an agent that should have been running against the dev DB was running against the production one.
  • Resource requirements from all experiments are provided

Edit | Attach | Watch | Print version | History: r9 < r8 < r7 < r6 < r5 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r9 - 2005-08-15 - unknown
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback